OpenStudy (anonymous):

circle centered at (2,3) and radius is square root 5 what is the equation?

OpenStudy (aakashsudhakar):

Equation for a circle is... (x - h)^2 + (y - k)^2 = r^2 where (h, k) are the coordinates of the center of the circle and r is the radius. Therefore, simply plug in values to get your proper equation! (h, k) --> (2, 3) r = √(5) (x - h)^2 + (y - k)^2 = r^2 (x - 2)^2 + (y - 3)^2 = 5 Thus, your final answer is \[(x - 2)^2 + (y - 3)^2 = 5\]

OpenStudy (anonymous):

Thanks,

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.